# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 3.2.2 Optimization Characteristics of Models
Optimization Analysis of Transformers under ICL Regimes. For the first aspect, Shen2025 theoretically investigate the training dynamics of a single-layer Transformer model for in-context classification tasks on Gaussian mixtures, demonstrating that under optimization via gradient descent, the model can converge to the global optimum at a linear rate. Zhang2024 study the training dynamics of a Transformer with a single linear attention layer during in-context learning for linear regression tasks, showing that the model can find the global minimum of the objective function. Chen2024unveiling use gradient flow to analyze how a simplified Transformer architecture with two attention layers performs ICL, revealing the collaborative mechanism of its components. Gong2025disentangling analyze the optimization dynamics of a single-layer Transformer with normalized ReLU self-attention under ICL mechanisms, indicating that smaller eigenvalues preserve basic knowledge, while larger eigenvalues of attention weights capture specialized knowledge. Zheng2024 examines whether autoregressively trained Transformers implement ICL by learning a meta-optimizer. They demonstrate that under specific conditions on the initial data distribution, the trained Transformer indeed learns to perform one-step gradient descent to solve ordinary least squares (OLS) problems in-context. Chen2024training prove that the training dynamics consist of three phases: warm-up, emergence, and convergence, with ICL capabilities rapidly emerging during the emergence phase. Huang2024 further extend previous analyses from linear attention to softmax attention, demonstrating that on balanced data, the model converges to near-zero prediction error through a two-phase process. On imbalanced data, the model exhibits “staged” convergence. Kim2024 theoretically study how Transformers with both MLP and attention layers learn nonlinear features in in-context learning (ICL). Under the assumption that the attention layers converge rapidly, the authors show that the infinite-dimensional loss landscape for MLP parameters exhibits a benign non-convex structure.
@@
Optimization Analysis of Transformers without ICL Regimes. The aforementioned studies mainly focus on demystifying the mechanism of ICL. Apart from them, there are also studies that directly study the training behaviors of transformers without ICL. Nichani2025 demonstrate that a single-layer Transformer with self-attention and MLP can achieve perfect prediction accuracy when the number of self-attention parameters or MLP parameters scales almost linearly with the number of facts. Tian2023 reveal that the self-attention mechanism exhibits a “scan and snap” dynamic: initially distributing attention uniformly across all tokens, it gradually focuses on “distinctive” tokens that are discriminative for predicting specific next tokens while reducing attention on “common” tokens that frequently appear across different next-token predictions. Ren2024 analyze the training dynamics of a single-layer Transformer on a synthetic dataset, showing that the optimization process consists of a “sample-intensive” stage and “sample-efficient” stage. Pan2025 propose a mathematical framework based on compression theory to explain the behavior of LLMs. They conceptualize LLM training as first learning and compressing common syntactic patterns, then progressively acquiring and storing knowledge from common to rare.
@@
Optimization Landscape. In addition to directly analyzing the optimization dynamics of transformers, some studies have also explored the model’s optimization process from the perspective of loss landscapes. Wen2025 propose the “River Valley Loss Landscape” hypothesis to analyze the effectiveness of the Warmup-Stable-Decay (WSD) learning rate schedule. Through theoretical analysis, they demonstrate that during the stable learning rate phase, a higher learning rate causes parameters to oscillate significantly between the “hillsides”, while also enabling faster progress along the direction of the “river” at the bottom. In the decay phase, the rapidly decreasing learning rate reduces oscillation amplitude, allowing parameters to move closer to the “river”, leading to a swift decrease in loss. Similarly, Liu2025, drawing from principles of classical thermodynamics, argue that the model training process can be decomposed into two dynamical stages: a fast dynamics phase characterized by rapid oscillations between hillsides, and a slow dynamics phase involving gradual drift along the river direction. Gong2025what further extend the “River Valley Loss Landscape” framework by proposing two types of river valleys: U-shaped and V-shaped valleys. U-shaped valleys are wide and flat, where optimization tends to stagnate. In contrast, V-shaped valleys feature narrow bottoms and steep sides, allowing parameters to “jump” between valley walls while progressing along the river direction.
@@ 3.2.3 Theoretical Design of Models
Unrolled optimization perspective. A mainstream and widely appreciated principle is to relate the forward pass of the stacked layers to an unrolled optimization process [ISTA, unroll-dictionary, hinton2022forward, monga2021algorithm, zhang2018ista, guo2023contranorm]. Given the input $\bm{z}$} and some model $f$ with $L$ layers, the core idea of the unrolled optimization is to interpret the layer-wise computation of the model as performing an iterative optimization on some latent objective function $F$, that is,
@@
where $f^{l}$ is the $l$-the layer and ${\bm{z}}^{l}$ is its corresponding output. In other words, each layer of the model can be viewed as a single step of an optimization algorithm seeking to minimize or optimize the underlying objective. Given the success of the Transformer architecture, it is natural for researchers to attempt to understand its structure from principled perspectives [whiteTF, whiteTF-JMLR, whiteTF2, IB-TF, attention-only-TF, yang2022transformers, ren2025transformers]. One prominent direction is to relate the Transformer’s design to an objective function associated with information compression. whiteTF showed that Transformer-like deep network layers can naturally be connected to an optimization process aimed at sparse rate reduction. More specifically, given the input data ${\bm{X}}\in{\mathbb{R}}^{d\times N}$, they denote $({\bm{U}}_{k})_{k=1}^{K}$ to be the set of bases of the mixture of low-dimensional $K$ Gaussian distributions. Then the objective function $F$ can be formalized as
@@
where $R$ and $R^{c}$ are estimates of lossy coding rates [ma2007segmentation, yu2020learning, redunet]. The objective aims to maximize the information gain for the final token representations by maximizing $\Delta R$ while promoting the sparsity by minimizing the $\ell^{0}$ norm. It is worth noting that the optimization of $\Delta R$ shares the same underlying inspiration as the design of ReduNet [redunet]. However, here the optimization of $R^{c}$ ultimately gives rise to the multi-head attention structure, whereas the remaining optimization of $R(\bm{Z})-\lambda\|\bm{Z}\|_{0}$ corresponds to a structure analogous to a feed-forward network (FFN) [ISTA].
In addition to the interpretation based on rate reduction, IB-TF approached the emerging visual grouping phenomenon observed in Vision Transformers from the perspective of the information bottleneck. They showed that the iterative solution to the information bottleneck objective can be expressed in the form of self-attention. attention-only-TF pointed out that compressing noisy token representations and the corresponding denoising operations can naturally give rise to the form of multi-head self-attention. Other works related the optimization objective $F$ to energy-based principles [HN-is-all-you-need, EnergyTF, hu2023sparse, wu2023stanhop, ren2025transformers, hu2025hyper].
Test-time Training Perspective. Although Transformers have achieved widespread success across tasks in different modalities, their quadratic complexity with respect to sequence length often becomes unacceptable under resource-constrained conditions. In addition to improving the efficiency of the Transformer architecture itself, researchers have also begun to focus on designing more efficient model architectures, among which an important line of work is called test-time training (regression) in TTT, yang2023gated, von2025mesanet, wang2025test, behrouz2024titans, behrouz2025s.
Generally, the design of this framework can be roughly divided into two stages [wang2025test]. The first stage uses a function $f_{t}$ to store memory in a regression manner at $t$-th test step, and the second stage uses this function for retrieval. Formally, similar to attention mechanisms, we transform the token $\bm{z}$} into the form of a query $\bm{q}$}, and convert the previous $t$ interacted tokens $\bm{x}$} into key-value pairs $({\bm{k}}_{1},{\bm{v}}_{1}),({\bm{k}}_{2},{\bm{v}}_{2}),\dots,({\bm{k}}_{t},{\bm{v}}_{t})$. Then, the output ${\bm{y}}_{t}$ at $t$-th step can be formalized as
@@
In fact, when we extend the above case using the kernel trick, where a kernel feature map $\phi$ is used to strengthen the representation of ${\bm{k}}_{t}$ and ${\bm{q}}_{t}$ [elu+1, performer]. This leads to the unnormalized softmax attention, which is also referred to as the dual model by ren2024towards. In addition, when we assign different weights in the linear-attention setting, this gives rise to the family of gated linear attentions such as RetNet [RetNet, yang2023gated, LRU, RWKV]. When the weight $\gamma_{i}$ depends on the input, it is often interpreted as a forget gate. Existing studies also show that state-space models can be viewed as a branch of gated linear attention [Mamba2, Mamba_thu_linear_attention, ren2024exploring].
Furthermore, when we consider online or streaming setting and apply different gradient-descent algorithms, we obtain the existing family of linear models known as fast weight programmers and online learners [schmidhuber1992learning, DeltaNet, longhorn, GatedDeltanet]. More specifically, if we perform single-example SGD at each time step and initialize $\bm{W}$} using the linear mapping obtained in the previous step, we can get the form of Delta Rule [DeltaNet], that is, ${\bm{W}}_{t}={\bm{W}}_{t-1}({\bm{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{T})+\beta_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{T}$ where $\beta_{t}$ is the learning rate at time $t$. Longhorn [longhorn] extends the above form by using the adaptive step sizes. In addition, Gated DeltaNet [GatedDeltanet] combines DeltaNet [DeltaNet] with the forget-gate mechanism of Mamba-2 [Mamba2], which can be viewed as adding an $F$-norm regularization on $\bm{W}$} to the original objective and performing single-example SGD.
@@ 3.3.1 Linear Models & No free Lunch
As discussed in Section 3.2.3, despite the remarkable performance of Transformers across a wide range of tasks, their quadratic computational cost remains a significant obstacle to broad deployment in real-world settings [transformer]. This has motivated a surge of interest in more efficient architectures whose computational and memory costs scale linearly with sequence length, including RetNet [RetNet], RWKV [RWKV], gated linear attention [yang2023gated], TTT [TTT], Mamba [Mamba, Mamba2], Longhorn [longhorn], gated DeltaNet [DeltaNet, GatedDeltanet]. These models are now widely recognized as belonging to the family of linear RNNs or as instances of the test-time training paradigm [yang2023gated, wang2025test]. However, the well-known “no free lunch” principle quickly comes into play: linear models often gain efficiency at the expense of representational power. Intuitively, as such models must compress past information into a fixed-size state without knowing what future inputs will be, two inherent difficulties arise. First, a constant-size state cannot scale with sequence length, causing substantial information loss on long inputs. Second, if future patterns deviate from the prior encoded in this compression rule, the compressed representation may completely fail. These limitations are reflected in recent theoretical findings. phonebook show that Transformers can copy sequences of exponential length, whereas fixed-state models are fundamentally limited by their finite memory. Similarly, Lvkaifeng_rnn_not_tf demonstrate that generalized RNNs even equipped with chain-of-thought reasoning cannot perform associative recall or other tasks requiring precise contextual retrieval unless they are augmented with retrieval-augmented generation (RAG) or followed by a Transformer layer.
@@
Even so, this does not imply that RNNs are necessarily weaker than Transformers. For example, bhattamishra2024separations prove that bounded Dyck languages can be recognized by constant-size RNNs, while a single-layer Transformer requires linear width. merrill_mamba further show that the expressive power of linear RNNs with diagonal transition matrices is comparable to that of Transformers (both lying within $TC^{0}$), yet allowing data-dependent non-diagonal transitions enables linear RNNs to surpass $TC^{0}$ class [merrill_mamba, grazzi2024unlocking, siems2025deltaproduct]. These observations point toward an appealing research direction: Hybrid architectures that combine linear models with Transformers. Lvkaifeng_rnn_not_tf theoretically show that simply introducing just one single Transformer layer into RNN is sufficient to enhance its in-context retrieval capability and close the representation gap with Transformers. Practical evidence also suggests that hybrid architectures, such as combining Mamba with Transformers, can achieve high efficiency while keeping comparable performance [Mamba_large_ex, Jamba, Zamba]. In addition, recent work has explored incorporating the Delta Rule into Transformers to further strengthen their expressive power [zhong2025understanding, deltaformer].
@@ 3.3.2 Recurrent Models & Looped Transformers
Beyond the pursuit of more efficient linear models, recurrent architectures have also begun to re-enter the spotlight in the community. This renewed interest is driven by several factors. On one hand, the emergence of chain-of-thought (CoT) reasoning has dramatically boosted models’ expressive power [CoT, hedi_CoT, malach2023auto, merrill2023expressive, li2024chain], prompting researchers to consider how such iterative reasoning capabilities might be implicitly baked into the model’s inductive bias [yu2025enhancing]. More broadly, a strengthened understanding of scaling laws has highlighted that performance gains come not only from scaling data and model size during training [scalinglaw1, scalinglaw2], but also from increasing test-time computation [test-time-scalinglaw], for example, by allowing the model to perform recurrent or iterative reasoning [geiping2025scaling, zhu2025scaling, wu2025parallel]. In fact, the study of recurrent or weight-tied architectures has a long and rich history, providing a strong foundation for these recent developments.
@@
As discussed in Section 3.2.1, dehghani2018universal introduced the Universal Transformer, which improves generalization by sharing parameters across layers and allowing the model to flexibly adjust its iterative depth. giannou2023looped further proposed treating Transformers as programmable computational units, where a fixed layer is repeatedly applied to execute instructions encoded in the input sequence. looptf-yangliu incorporated the looping paradigm directly into the Transformer’s iterative computation process, enabling the model to more effectively learn tasks that require internal learning algorithms. gatmiry2024can studied whether looped Transformers can implement multi-step gradient descent in an in-context learning setting. looptf-fan, looptf-hedi demonstrated that looped Transformers achieve substantially better length generalization compared to fixed-depth Transformers. saunshi2025reasoning demonstrated that many reasoning problems require greater depth rather than more parameters, and that looped models can achieve more effective reasoning while using significantly fewer parameters. Collectively, these studies highlight the advantages of recurrence primarily through theoretical analyses or small-scale experiments. More recently, geiping2025scaling used recurrence as a prior for implicit reasoning in latent space, scaling the model to 3.5B parameters and showing performance competitive with non-looped models tens of billions of parameters in size. Similarly, loopTF-scaling introduced Ouro, a family of pre-trained looped language models scaling up to 2.6B parameters and trained on 7.7T tokens. loopTF-parallel proposed the Parallel Loop Transformer (PLT) architecture to improve computational efficiency when leveraging recurrence. MoR developed Mixture-of-Recursions (MoR), which combines parameter sharing with adaptive computation to unlock stronger model performance.

### v2

@@ 3.2.2 Optimization Characteristics of Models
Optimization Analysis of Transformers under ICL Regimes. For the first aspect, Shen et al. (2025) theoretically investigate the training dynamics of a single-layer Transformer model for in-context classification tasks on Gaussian mixtures, demonstrating that under optimization via gradient descent, the model can converge to the global optimum at a linear rate. Zhang et al. (2024b) study the training dynamics of a Transformer with a single linear attention layer during in-context learning for linear regression tasks, showing that the model can find the global minimum of the objective function. Chen et al. (2024e) use gradient flow to analyze how a simplified Transformer architecture with two attention layers performs ICL, revealing the collaborative mechanism of its components. Gong et al. (2025a) analyze the optimization dynamics of a single-layer Transformer with normalized ReLU self-attention under ICL mechanisms, indicating that smaller eigenvalues preserve basic knowledge, while larger eigenvalues of attention weights capture specialized knowledge. Zheng et al. (2024) examines whether autoregressively trained Transformers implement ICL by learning a meta-optimizer. They demonstrate that under specific conditions on the initial data distribution, the trained Transformer indeed learns to perform one-step gradient descent to solve ordinary least squares (OLS) problems in-context. Chen et al. (2024d) prove that the training dynamics consist of three phases: warm-up, emergence, and convergence, with ICL capabilities rapidly emerging during the emergence phase. Huang et al. (2024c) further extend previous analyses from linear attention to softmax attention, demonstrating that on balanced data, the model converges to near-zero prediction error through a two-phase process. On imbalanced data, the model exhibits “staged” convergence. Kim and Suzuki (2024) theoretically study how Transformers with both MLP and attention layers learn nonlinear features in in-context learning (ICL). Under the assumption that the attention layers converge rapidly, the authors show that the infinite-dimensional loss landscape for MLP parameters exhibits a benign non-convex structure.
@@
Optimization Analysis of Transformers without ICL Regimes. The aforementioned studies mainly focus on demystifying the mechanism of ICL. Apart from them, there are also studies that directly study the training behaviors of transformers without ICL. Nichani et al. (2025) demonstrate that a single-layer Transformer with self-attention and MLP can achieve perfect prediction accuracy when the number of self-attention parameters or MLP parameters scales almost linearly with the number of facts. Tian et al. (2023) reveal that the self-attention mechanism exhibits a “scan and snap” dynamic: initially distributing attention uniformly across all tokens, it gradually focuses on “distinctive” tokens that are discriminative for predicting specific next tokens while reducing attention on “common” tokens that frequently appear across different next-token predictions. Ren et al. (2024b) analyze the training dynamics of a single-layer Transformer on a synthetic dataset, showing that the optimization process consists of a “sample-intensive” stage and “sample-efficient” stage. Pan et al. (2025b) propose a mathematical framework based on compression theory to explain the behavior of LLMs. They conceptualize LLM training as first learning and compressing common syntactic patterns, then progressively acquiring and storing knowledge from common to rare.
@@
Optimization Landscape. In addition to directly analyzing the optimization dynamics of transformers, some studies have also explored the model’s optimization process from the perspective of loss landscapes. Wen et al. (2025) propose the “River Valley Loss Landscape” hypothesis to analyze the effectiveness of the Warmup-Stable-Decay (WSD) learning rate schedule. Through theoretical analysis, they demonstrate that during the stable learning rate phase, a higher learning rate causes parameters to oscillate significantly between the “hillsides”, while also enabling faster progress along the direction of the “river” at the bottom. In the decay phase, the rapidly decreasing learning rate reduces oscillation amplitude, allowing parameters to move closer to the “river”, leading to a swift decrease in loss. Similarly, Liu et al. (2025i), drawing from principles of classical thermodynamics, argue that the model training process can be decomposed into two dynamical stages: a fast dynamics phase characterized by rapid oscillations between hillsides, and a slow dynamics phase involving gradual drift along the river direction. Gong et al. (2025b) further extend the “River Valley Loss Landscape” framework by proposing two types of river valleys: U-shaped and V-shaped valleys. U-shaped valleys are wide and flat, where optimization tends to stagnate. In contrast, V-shaped valleys feature narrow bottoms and steep sides, allowing parameters to “jump” between valley walls while progressing along the river direction.
@@ 3.2.3 Theoretical Design of Models
Unrolled optimization perspective. A mainstream and widely appreciated principle is to relate the forward pass of the stacked layers to an unrolled optimization process (Gregor and LeCun, 2010; Tolooshams and Ba, 2021; Hinton, 2022; Monga et al., 2021; Zhang and Ghanem, 2018; Guo et al., 2023c). Given the input $\bm{z}$} and some model $f$ with $L$ layers, the core idea of the unrolled optimization is to interpret the layer-wise computation of the model as performing an iterative optimization on some latent objective function $F$, that is,
@@
where $f^{l}$ is the $l$-the layer and ${\bm{z}}^{l}$ is its corresponding output. In other words, each layer of the model can be viewed as a single step of an optimization algorithm seeking to minimize or optimize the underlying objective. Given the success of the Transformer architecture, it is natural for researchers to attempt to understand its structure from principled perspectives (Yu et al., 2023a; 2024; b; Zhou et al., 2022; Wang et al., 2025b; Yang et al., 2022; Ren et al., 2025). One prominent direction is to relate the Transformer’s design to an objective function associated with information compression. Yu et al. (2023a) show that Transformer-like deep network layers can naturally be connected to an optimization process aimed at sparse rate reduction. More specifically, given the input data ${\bm{X}}\in{\mathbb{R}}^{d\times N}$, they denote $({\bm{U}}_{k})_{k=1}^{K}$ to be the set of bases of the mixture of low-dimensional $K$ Gaussian distributions. Then the objective function $F$ can be formalized as
@@
where $R$ and $R^{c}$ are estimates of lossy coding rates (Ma et al., 2007; Yu et al., 2020; Chan et al., 2022). The objective aims to maximize the information gain for the final token representations by maximizing $\Delta R$ while promoting the sparsity by minimizing the $\ell^{0}$ norm. It is worth noting that the optimization of $\Delta R$ shares the same underlying inspiration as the design of ReduNet (Chan et al., 2022). However, here the optimization of $R^{c}$ ultimately gives rise to the multi-head attention structure, whereas the remaining optimization of $R(\bm{Z})-\lambda\|\bm{Z}\|_{0}$ corresponds to a structure analogous to a feed-forward network (FFN) (Gregor and LeCun, 2010).
In addition to the interpretation based on rate reduction, Zhou et al. (2022) approach the emerging visual grouping phenomenon observed in Vision Transformers from the perspective of the information bottleneck. They showed that the iterative solution to the information bottleneck objective can be expressed in the form of self-attention. Wang et al. (2025b) point out that compressing noisy token representations and the corresponding denoising operations can naturally give rise to the form of multi-head self-attention. Other works relate the optimization objective $F$ to energy-based principles (Ramsauer et al., 2020; Hoover et al., 2023; Hu et al., 2023a; Wu et al., 2023; Ren et al., 2025; Hu et al., 2025).
Test-time Training Perspective. Although Transformers have achieved widespread success across tasks in different modalities, their quadratic complexity with respect to sequence length often becomes unacceptable under resource-constrained conditions. In addition to improving the efficiency of the Transformer architecture itself, researchers have also begun to focus on designing more efficient model architectures, among which an important line of work is called test-time training (regression) in Sun et al. (2024); Yang et al. (2023c); von Oswald et al. (2025); Wang et al. (2025a); Behrouz et al. (2024; 2025).
Generally, the design of this framework can be roughly divided into two stages (Wang et al., 2025a). The first stage uses a function $f_{t}$ to store memory in a regression manner at $t$-th test step, and the second stage uses this function for retrieval. Formally, similar to attention mechanisms, we transform the token $\bm{z}$} into the form of a query $\bm{q}$}, and convert the previous $t$ interacted tokens $\bm{x}$} into key-value pairs $({\bm{k}}_{1},{\bm{v}}_{1}),({\bm{k}}_{2},{\bm{v}}_{2}),\dots,({\bm{k}}_{t},{\bm{v}}_{t})$. Then, the output ${\bm{y}}_{t}$ at $t$-th step can be formalized as
@@
In fact, when we extend the above case using the kernel trick, where a kernel feature map $\phi$ is used to strengthen the representation of ${\bm{k}}_{t}$ and ${\bm{q}}_{t}$ (Katharopoulos et al., 2020; Choromanski et al., 2020). This leads to the unnormalized softmax attention, which is also referred to as the dual model by Ren and Liu (2024). In addition, when we assign different weights in the linear-attention setting, this gives rise to the family of gated linear attentions such as RetNet (Sun et al., 2023; Yang et al., 2023c; Orvieto et al., 2023; Peng et al., 2023). When the weight $\gamma_{i}$ depends on the input, it is often interpreted as a forget gate. Existing studies also show that state-space models can be viewed as a branch of gated linear attention (Dao and Gu, 2024; Han et al., 2024; Ren et al., 2024a).
Furthermore, when we consider online or streaming setting and apply different gradient-descent algorithms, we obtain the existing family of linear models known as fast weight programmers and online learners (Schmidhuber, 1992; Yang et al., 2024b; Liu et al., 2024a; Yang et al., 2024c). More specifically, if we perform single-example SGD at each time step and initialize $\bm{W}$} using the linear mapping obtained in the previous step, we can get the form of Delta Rule (Yang et al., 2024b), that is, ${\bm{W}}_{t}={\bm{W}}_{t-1}({\bm{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{T})+\beta_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{T}$ where $\beta_{t}$ is the learning rate at time $t$. Longhorn (Liu et al., 2024a) extends the above form by using the adaptive step sizes. In addition, Gated DeltaNet (Yang et al., 2024c) combines DeltaNet (Yang et al., 2024b) with the forget-gate mechanism of Mamba-2 (Dao and Gu, 2024), which can be viewed as adding an $F$-norm regularization on $\bm{W}$} to the original objective and performing single-example SGD.
@@ 3.3.1 Linear Models & No free Lunch
As discussed in Section 3.2.3, despite the remarkable performance of Transformers across a wide range of tasks, their quadratic computational cost remains a significant obstacle to broad deployment in real-world settings (Vaswani et al., 2017a). This has motivated a surge of interest in more efficient architectures whose computational and memory costs scale linearly with sequence length, including RetNet (Sun et al., 2023), RWKV (Peng et al., 2023), gated linear attention (Yang et al., 2023c), TTT(Sun et al., 2024), Mamba(Gu and Dao, 2023; Dao and Gu, 2024), Longhorn (Liu et al., 2024a), gated DeltaNet (Yang et al., 2024b; c). These models are now widely recognized as belonging to the family of linear RNNs or as instances of the test-time training paradigm (Yang et al., 2023c; Wang et al., 2025a). However, the well-known “no free lunch" principle quickly comes into play: linear models often gain efficiency at the expense of representational power. Intuitively, as such models must compress past information into a fixed-size state without knowing what future inputs will be, two inherent difficulties arise. First, a constant-size state cannot scale with sequence length, causing substantial information loss on long inputs. Second, if future patterns deviate from the prior encoded in this compression rule, the compressed representation may completely fail. These limitations are reflected in recent theoretical findings. Jelassi et al. (2024) show that Transformers can copy sequences of exponential length, whereas fixed-state models are fundamentally limited by their finite memory. Similarly, Wen et al. (2024) demonstrate that generalized RNNs even equipped with chain-of-thought reasoning cannot perform associative recall or other tasks requiring precise contextual retrieval unless they are augmented with retrieval-augmented generation (RAG) or followed by a Transformer layer.
@@
Even so, this does not imply that RNNs are necessarily weaker than Transformers. For example, Bhattamishra et al. (2024) prove that bounded Dyck languages can be recognized by constant-size RNNs, while a single-layer Transformer requires linear width. Merrill et al. (2024) further show that the expressive power of linear RNNs with diagonal transition matrices is comparable to that of Transformers (both lying within $TC^{0}$), yet allowing data-dependent non-diagonal transitions enables linear RNNs to surpass $TC^{0}$ class (Merrill et al., 2024; Grazzi et al., 2024; Siems et al., 2025). These observations point toward an appealing research direction: Hybrid architectures that combine linear models with Transformers. Wen et al. (2024) theoretically show that simply introducing just one single Transformer layer into RNN is sufficient to enhance its in-context retrieval capability and close the representation gap with Transformers. Practical evidence also suggests that hybrid architectures, such as combining Mamba with Transformers, can achieve high efficiency while keeping comparable performance (Waleffe et al., 2024; Lieber et al., 2024; Glorioso et al., 2024). In addition, recent work has explored incorporating the Delta Rule into Transformers to further strengthen their expressive power (Zhong et al., 2025b; Xu et al., 2025b).
@@ 3.3.2 Recurrent Models & Looped Transformers
Beyond the pursuit of more efficient linear models, recurrent architectures have also begun to re-enter the spotlight in the community. This renewed interest is driven by several factors. On one hand, the emergence of chain-of-thought (CoT) reasoning has dramatically boosted models’ expressive power (Wei et al., 2022c; Feng et al., 2024a; Malach, 2023; Merrill and Sabharwal, 2023a; Li et al., 2024b), prompting researchers to consider how such iterative reasoning capabilities might be implicitly baked into the model’s inductive bias (Yu et al., 2025b). More broadly, a strengthened understanding of scaling laws has highlighted that performance gains come not only from scaling data and model size during training (Kaplan et al., 2020b; Hoffmann et al., 2022a), but also from increasing test-time computation (Snell et al., 2024b), for example, by allowing the model to perform recurrent or iterative reasoning (Geiping et al., 2025; Zhu et al., 2025e; Wu et al., 2025b). In fact, the study of recurrent or weight-tied architectures has a long and rich history, providing a strong foundation for these recent developments.
@@
As discussed in Section 3.2.1, Dehghani et al. (2018) introduce the Universal Transformer, which improves generalization by sharing parameters across layers and allowing the model to flexibly adjust its iterative depth. Giannou et al. (2023b) further propose treating Transformers as programmable computational units, where a fixed layer is repeatedly applied to execute instructions encoded in the input sequence. Yang et al. (2023b) incorporate the looping paradigm directly into the Transformer’s iterative computation process, enabling the model to more effectively learn tasks that require internal learning algorithms. Gatmiry et al. (2024) study whether looped Transformers can implement multi-step gradient descent in an in-context learning setting. Fan et al. (2024b); Yu et al. (2025a) demonstrate that looped Transformers achieve substantially better length generalization compared to fixed-depth Transformers. Saunshi et al. (2025b) demonstrate that many reasoning problems require greater depth rather than more parameters, and that looped models can achieve more effective reasoning while using significantly fewer parameters. Collectively, these studies highlight the advantages of recurrence primarily through theoretical analyses or small-scale experiments. More recently, Geiping et al. (2025) use recurrence as a prior for implicit reasoning in latent space, scaling the model to 3.5B parameters and showing performance competitive with non-looped models tens of billions of parameters in size. Similarly, Zhu et al. (2025d) introduce Ouro, a family of pre-trained looped language models scaling up to 2.6B parameters and trained on 7.7T tokens. Wu et al. (2025a) propose the Parallel Loop Transformer (PLT) architecture to improve computational efficiency when leveraging recurrence. Bae et al. (2025) develope Mixture-of-Recursions (MoR), which combines parameter sharing with adaptive computation to unlock stronger model performance.